A boat traveled 112miles downstream took 7hrs. The trip back took 14hrs. What is the speed of the boat in still water? What is the speed of the current?
Answer by flame8855(424)

let x be the speed of boat in still water
let c be the speed of current
downstream : speed = x + c
x+c= 112/7 = 16
upstream : speed = x-c
x-c= 112/14 = 8
solve
x= 12 mph speed of boat in still water
c = 4 mph speed of current